OpenStudy (anonymous):

Simplify. 26 + (–3) ∙ (–8) plz plz help

OpenStudy (alexandervonhumboldt2):

first multiply: (-3)*(-8)=3*8=24 -*-=+. now add: 26+24=50

OpenStudy (alexandervonhumboldt2):

Use PEDMAS

OpenStudy (anonymous):

ohh ok thanks soo much